A heart pacemaker fires 89 times a minute, each time a 27.0 nF capacitor is charged (by a battery in series with a resistor) to 0.632 of its full voltage. What is the value of the resistance (in Ω)?
Answer #1

Solution :-

Time for charging the capacitor = (60/89) s = 0.674 s

In this time interval capacitor is charged to 0.632 times of its full voltage, which indicates that this is the value of time constant of the series RC circuit. By relation we know that,

R = (0.674)÷(27×10^-9)

= 24.96×10^6 ohm